Project

General

Profile

Actions

2017-11-30 physics-shifter

Start Date: 2017-11-30
End Date: 2017-12-05
Start Run: 10313
End Run: 10431
Physics shifter: Chris Rogers

Summary

The aim was to get the empty data for corresponding LiH full settings (solenoid mode)

  • 30th Nov - 1st Dec: 2017-02-6
  • 1st Dec - 2nd Dec: 2016-04-1.2
  • 2nd Dec - 3rd Dec: 2016-04-1.7
  • 3rd Dec - 4th Dec: 2016-04-2.4a
  • 4th Dec - SSD quenched

Note that KL was fixed around 30th Nov.

Elog, CDB

Fine

Online reco

Fine

Analysis

30th November

Checked:
  • TOF1 dt is 50 ps, still not great; TOF0 and TOF2 slab conversion to space point efficiency ~ 97-98 %
  • TKU, TKD has no obvious dead area
  • Amplitude change looks good

1st December

Checked:
  • TOF1 dt is 50 ps, still not great; TOF0 and TOF2 slab conversion to space point efficiency ~ 97-98 %
  • TKU, TKD has no obvious dead area; chi2 looks reasonable
  • Amplitude change looks good

I also did some beamline optimisation with Paolo; results here

2nd December - 200 MeV/c channel

Checked as per above; all looks good.

3rd December - 240 MeV/c channel

Checked as per above; all looks good. Noted that chi2 distribution appears to have less tail for 240 MeV/c data...

4th December

The magnets quenched - straight tracks data overnight

Updated by Rogers, Chris almost 5 years ago · 12 revisions